xx InitBase

InitBase

Erik Schnetter <schnetter@cct.lsu.edu>

July 06 2007

Abstract

Thorn InitBase specifis how initial data are to be set up. It does not set up any initial data by itself, nor does it contain any routines which are to be called. It is merely a convenient repository remembering how initial data are to be set up, so that other thorns can check their actions against this thorn.

There are several possibilities:

The initial data thorn sets up data on one time level, while other time levels are scratch space. The time evolution method must start up from a single time level. (This is the default.)

The initial data thorn sets up data on exactly one time level, and is called once for each active time level. (This means that the initial data thorn can only access the current time level.)

The initial data thorn sets up data on exactly two time levels, and is called once for each active time level. (This means that the initial data thorn can only access the current and the first past time level.)

The initial data thorn sets up data on all active time levels. (This makes it necessary that the initial data thorn checks the number of active time levels.)

This thorn does not set up any initial data by itself. It only selects the mechanism; it is the responsibility of other thorns to follow this specification. If all other thorns follow this specification, then they are guaranteed to interact nicely. If other thorns ignore this specification, then end-users may be surprised and may have difficulties setting up parameter files which work correctly.

1 Parameters

11initialxdataxsetupxmethod 10Scope: restrictedx01  KEYWORD

11Description: Procedure for setting up initial data

11 Range 10 x01 Default: initxsomexlevels
11initxsomexlevels  10Set up at least one time level; other time levels are scratch space
11initxsinglexlevel  10Set up exactly one time level; other time levels are not accessed
11initxtwoxlevels  10Set up exactly two time levels; other time levels are not accessed
11initxallxlevels  10Set up all active time levels

11

2 Interfaces

General Implements: initbase

3 Schedule This section lists all the variables which are assigned storage by thorn CactusBase/InitBase. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE